OpenStudy (anonymous):

find solutions for x: sqrt{x^2-11=5}

OpenStudy (anonymous):

\[\sqrt{x^2-11=5}\]

OpenStudy (anonymous):

you mean \[\sqrt{x^2-11} =5\] most likely

OpenStudy (anonymous):

sqaure both sides x^2 -11 = 25 x^2 = 36 x= +- 6

OpenStudy (anonymous):

BUT

OpenStudy (anonymous):

it has no solutions

OpenStudy (anonymous):

because if you put either +6 or -6 into the LHS you end up needing to take the sqrt of a negative

OpenStudy (anonymous):

so , has no real solutions

OpenStudy (anonymous):

you need to check your answers by subbing in if you ever have a sqaure root or a logarithm or divide by a variable